# Invariant subspaces of Brownian shifts on vector-valued Hardy spaces

### v1

@@ 1. Introduction
where $\varphi$ is an inner function [5]. Brownian shifts were introduced by Agler and Stankus on $H^{2}(\mathbb{T})\oplus\mathbb{C}$ [3, Definition 5.5] in the context of $m$-isometries, and, along the lines of Beurling, they characterized the invariant subspaces of such operators. The Brownian shift of covariance $\sigma>0$ and angle $\theta\in[0,2\pi)$ is the bounded linear operator $B_{\sigma,e^{i\theta}}:H^{2}(\mathbb{T})\oplus\mathbb{C}\to H^{2}(\mathbb{T})\oplus\mathbb{C}$, defined by
@@
On the other hand, we recall that shifts on vector-valued Hardy spaces are all examples of pure isometries. Given a Hilbert space $E$ (all Hilbert spaces in this paper are separable and over $\mathbb{C}$), denote by $H^{2}_{E}(\mathbb{T})$ the $E$-valued Hardy space over $\mathbb{T}$. By $S_{E}$, we refer to the shift operator on $H^{2}_{E}(\mathbb{T})$. When $E=\mathbb{C}$, we simply write $H^{2}_{\mathbb{C}}(\mathbb{T})$ as $H^{2}(\mathbb{T})$, and $S_{\mathbb{C}}$ as $S$. In this case, the invariant subspaces of $S_{E}$ are parameterized by operator-valued inner functions—a classical result known as the Beurling–Lax–Halmos theorem (cf. [11, Theorem 2.1, p. 239]). This result lies within the framework of the invariant subspace problem—where, instead of investigating the existence of invariant subspaces for general operators, one fixes a natural operator and seeks a complete description of the lattice of its invariant subspaces.
In this paper, we introduce Brownian shifts on vector-valued Hardy spaces and compute the lattice of invariant subspaces for these operators. This work is inspired both by the classification of invariant subspaces by Agler and Stankus and by the Beurling–Lax–Halmos theorem for shift-invariant subspaces on vector-valued Hardy spaces. In particular, our result yields the lattice of invariant subspaces for the Brownian shifts $B_{\sigma,e^{i\theta}}$ studied by Agler and Stankus and provides a new proof of their structure.
@@
This is a particular class of Brownian unitaries with positive covariance, as introduced by Agler and Stankus. Clearly, in the scalar case $E=\mathbb{C}$, we have $B^{\mathbb{C}}_{\sigma,e^{i\theta}}=B_{\sigma,e^{i\theta}}$. We often refer to $B_{\sigma,e^{i\theta}}$ as a Brownian shift on $H^{2}(\mathbb{T})$. Throughout the paper, $E$ will denote an arbitrary but fixed Hilbert space, with the possibility that $E=\mathbb{C}$. One additional motivation for introducing Brownian shifts on vector-valued Hardy spaces is that they are unitarily equivalent to the Brownian shifts on $H^{2}(\mathbb{T})$ tensored with identity operators. That is,
@@
We refer to Theorem 4.2 for the scalar version of the above result (also see [8, Theorem 2.1]). We finally remark that Brownian shifts are integral components of 222-isometries, which play an important role in understanding the general structure of linear operators (cf. [4, 7, 16]).
In the context of function theory, linear operators, and the prediction problem for 222-stationary processes, we refer the reader to the concluding part of the paper by Agler and Stankus [2]. For the prediction of stationary processes, we recommend the excellent monograph [6], particularly Chapter 2. See also [15, 17] for related work and possible directions for future connections.
The remainder of the paper is organized as follows. In Section 2, we describe the invariant subspaces of Brownian shifts. Section 3 determines when such pairs of invariant subspaces are unitarily equivalent. Section 4 presents the results obtained this far in the context of the scalar-valued Hardy space. In particular, it highlights the unitary equivalence result for Brownian shifts on $H^{2}(\mathbb{T})$. Section 5 presents the structure of reducing subspaces for $B_{\sigma,e^{i\theta}}^{E}$.
@@ 3. Equivalent invariant subspaces
$=U\left(\begin{bmatrix}\Phi_{1}zh\\ 0\end{bmatrix}+e^{i\theta}\begin{bmatrix}g_{1}\\ y_{1}\end{bmatrix}+\begin{bmatrix}\Phi_{1}x_{1}^{\prime}\\ 0\end{bmatrix}\right)$
$=U\left(\begin{bmatrix}z\Phi_{1}h\\ 0\end{bmatrix}+\begin{bmatrix}zg_{1}+\sigma_{1}y_{1}\\ e^{i\theta}y_{1}\end{bmatrix}\right)$
$=UB^{E}_{\sigma_{1},e^{i\theta}}\left(\begin{bmatrix}\Phi_{1}h\\ 0\end{bmatrix}+\begin{bmatrix}g_{1}\\ y_{1}\end{bmatrix}\right),$

### v2

@@ 1. Introduction
where $\varphi$ is an inner function. This is the classical result of Beurling [5]. Continuing with $S$ on $H^{2}(\mathbb{T})$, we now turn to Brownian shifts. These were introduced by Agler and Stankus on $H^{2}(\mathbb{T})\oplus\mathbb{C}$ [3, Definition 5.5] in the context of $m$-isometries, and, along the lines of Beurling, they characterized the invariant subspaces of such operators. The Brownian shift of covariance $\sigma>0$ and angle $\theta\in[0,2\pi)$ is the bounded linear operator $B_{\sigma,e^{i\theta}}:H^{2}(\mathbb{T})\oplus\mathbb{C}\to H^{2}(\mathbb{T})\oplus\mathbb{C}$, defined by
@@
On the other hand, we recall that shifts on vector-valued Hardy spaces are all examples of pure isometries. Given a Hilbert space $E$, denote by $H^{2}_{E}(\mathbb{T})$ the $E$-valued Hardy space over $\mathbb{T}$. By $S_{E}$, we refer to the shift operator on $H^{2}_{E}(\mathbb{T})$. When $E=\mathbb{C}$, we simply write $H^{2}_{\mathbb{C}}(\mathbb{T})$ as $H^{2}(\mathbb{T})$, and $S_{\mathbb{C}}$ as $S$. In this case, the invariant subspaces of $S_{E}$ are parameterized by operator-valued inner functions—a classical result known as the Beurling–Lax–Halmos theorem (cf. [11, Theorem 2.1, p. 239]). This result lies within the framework of the invariant subspace problem—where, instead of investigating the existence of invariant subspaces for general operators, one fixes a natural operator and seeks a complete description of the lattice of its invariant subspaces.
In this paper, we introduce Brownian shifts on vector-valued Hardy spaces and compute the lattice of invariant subspaces for these operators. This work is inspired both by the classification of invariant subspaces by Agler and Stankus and by the Beurling–Lax–Halmos theorem for shift-invariant subspaces on vector-valued Hardy spaces. In particular, our result yields the lattices of invariant subspaces for the Brownian shifts $B_{\sigma,e^{i\theta}}$ studied by Agler and Stankus and provides a new proof of their structure.
@@
This is a particular class of Brownian unitaries with positive covariance, as introduced by Agler and Stankus [3]. Clearly, in the scalar case $E=\mathbb{C}$, we have $B^{\mathbb{C}}_{\sigma,e^{i\theta}}=B_{\sigma,e^{i\theta}}$. We often refer to $B_{\sigma,e^{i\theta}}$ as a Brownian shift on $H^{2}(\mathbb{T})$. Throughout the paper, $E$ will denote a fixed, but otherwise arbitrary, separable Hilbert space over $\mathbb{C}$, with the possibility that $E=\mathbb{C}$. One additional motivation for introducing Brownian shifts on vector-valued Hardy spaces is that they are unitarily equivalent to the Brownian shifts on $H^{2}(\mathbb{T})$ tensored with identity operators. That is,
@@
We refer to Theorem 4.2 for the scalar version of the above result (also see [8, Theorem 2.1]).
A fundamental problem in Hilbert function space theory is the construction of new operators from existing ones. In particular, a natural question is whether the class of operators obtained by restricting natural operators to their invariant subspaces yields new operators (see the introductory part of Section 6 for more details).
In this context, Brownian shifts on $H^{2}_{E}(\mathbb{T})\oplus E$ exhibit mixed behavior, suggesting the potential for even richer structural properties. More specifically, in the final section, we apply our main results to show that there exists an invariant subspace of a Brownian shift such that the corresponding restriction operator is unitarily equivalent to another Brownian shift. At the same time, we also provide examples of invariant subspaces for which the restriction operator is not unitarily equivalent to any Brownian shift. These examples indicate that Brownian shifts exhibit behavior intermediate between that of shifts on the Hardy space—where restrictions to invariant subspaces remain unitarily equivalent shifts—and those on the Bergman and Dirichlet spaces, where such restrictions typically produce unitarily inequivalent operators.
We finally remark that Brownian shifts are integral components of $2$-isometries, which play an important role in understanding the general structure of linear operators (cf. [4, 7, 16]). In the context of function theory, linear operators, and the prediction problem for $2$-stationary processes, we refer the reader to the concluding part of the paper by Agler and Stankus [2]. For the prediction of stationary processes, we recommend the excellent monograph [6], particularly Chapter 2. See also [15, 17] for related work and possible directions for future connections.
The remainder of the paper is organized as follows. In Section 2, we describe the invariant subspaces of Brownian shifts. Section 3 determines when such pairs of invariant subspaces are unitarily equivalent. Section 4 presents the results obtained this far in the context of the scalar-valued Hardy space. In particular, it highlights the unitary equivalence result for Brownian shifts on $H^{2}(\mathbb{T})$. Section 5 presents the structure of reducing subspaces for $B_{\sigma,e^{i\theta}}^{E}$. The final section, Section 6, consists of concrete examples illustrating main results of this paper.
@@ 3. Equivalent invariant subspaces
[…]
@@ 5. Reducing subspaces
6. Examples
Let $\mathcal{M}\subseteq H^{2}_{E}(\mathbb{T})$ be a nonzero closed invariant subspace of the unilateral shift operator $S_{E}$. Then the restriction ${S_{E}}|_{\mathcal{M}}$ is itself a shift operator. In other words, up to unitary equivalence, the restriction of a shift operator to an invariant subspace is again a shift operator. Thus, from the perspective of operator theory, invariant subspaces of shift operators do not yield fundamentally new operators. The Brownian shift on $H^{2}(\mathbb{T})$ exhibits a similar property: up to unitary equivalence, the restriction of a Brownian shift on $H^{2}(\mathbb{T})$ to a Type II invariant subspace yields another Brownian shift on $H^{2}(\mathbb{T})$ [3, Proposition 5.75] (see also [8]).
At the other extreme, the situation is different for the Bergman and Dirichlet shifts on the unit disc. A rigidity result of Richter shows that the restrictions of the Bergman or Dirichlet shifts to their invariant subspaces always produce unitarily non-equivalent operators, that is, each invariant subspace gives rise to a distinct operator [20] (also see [9, 10]).
In this section, we present two concrete examples that illustrate the nuanced behavior of Brownian shifts. In the first example, we show that the restriction of a Brownian shift to an invariant subspace does not necessarily yield another Brownian shift:
Example 6.1.
Consider a Brownian shift $B_{\sigma,e^{i\theta}}^{E}$ acting on $H^{2}_{E}(\mathbb{T})\oplus E$ for some Hilbert space $E$ with $\dim(E)\geq 2$. Pick a nonzero closed subspace $E_{1}$ of $E$ such that
$\dim E_{1}\neq\dim E.$
By the upper triangular representation of Brownian shifts, it follows that $H^{2}_{E}(\mathbb{T})\oplus E_{1}$ is a Type II invariant subspace for $B_{\sigma,e^{i\theta}}^{E}$. If possible, suppose there exist $\sigma^{\prime}>0$, $\theta^{\prime}\in[0,2\pi)$ and a Hilbert space $G$ such that
$B_{\sigma,e^{i\theta}}^{E}\big{|}_{H^{2}_{E}(\mathbb{T})\oplus E_{1}}\cong B_{\sigma^{\prime},e^{i\theta^{\prime}}}^{G}.$ (6.1)
Consider the Brownian shift $B_{\sigma^{\prime},e^{i\theta^{\prime}}}^{K}$ on $H^{2}_{K}(\mathbb{T})\oplus K$, where
$K=E\oplus G.$
Clearly, there exists a unitary $U_{1}:H^{2}_{E}(\mathbb{T})\oplus E_{1}\rightarrow H^{2}_{E\oplus\{0\}}(\mathbb{T})\oplus(E_{1}\oplus\{0\})$ such that
$U_{1}B_{\sigma,e^{i\theta}}^{E}\big{|}_{H^{2}_{E}(\mathbb{T})\oplus E_{1}}=B_{\sigma,e^{i\theta}}^{K}\big{|}_{H^{2}_{E\oplus\{0\}}(\mathbb{T})\oplus(E_{1}\oplus\{0\})}U_{1}.$
Similarly, there is a unitary $U_{2}:H^{2}_{G}(\mathbb{T})\oplus G\rightarrow H^{2}_{\{0\}\oplus G}(\mathbb{T})\oplus(\{0\}\oplus G)$ such that
$U_{2}B_{\sigma^{\prime},e^{i\theta^{\prime}}}^{G}=B_{\sigma^{\prime},e^{i\theta^{\prime}}}^{K}\big{|}_{H^{2}_{\{0\}\oplus G}(\mathbb{T})\oplus(\{0\}\oplus G)}U_{2}.$
As a result, (6.1) implies
$B_{\sigma,e^{i\theta}}^{K}\big{|}_{H^{2}_{E\oplus\{0\}}(\mathbb{T})\oplus(E_{1}\oplus\{0\})}\cong B_{\sigma^{\prime},e^{i\theta^{\prime}}}^{K}\big{|}_{H^{2}_{\{0\}\oplus G}(\mathbb{T})\oplus(\{0\}\oplus G)}.$ (6.2)
On the other hand, observe that
$H^{2}_{E\oplus\{0\}}(\mathbb{T})\oplus(E_{1}\oplus\{0\})=\left(H^{2}_{E\oplus\{0\}}(\mathbb{T})\oplus\{0\}\right)\oplus\left(\{0\}\oplus(E_{1}\oplus\{0\})\right),$
and given any element $\begin{bmatrix}0\\ e_{1}\end{bmatrix}$, with $e_{1}\in E_{1}\oplus\{0\}$, we have
$0=\frac{\sigma e_{1}-\sigma e_{1}}{z-e^{i\theta}}.$
Therefore, it becomes apparent that the subspace $H^{2}_{E\oplus\{0\}}(\mathbb{T})\oplus(E_{1}\oplus\{0\})$ of $H^{2}_{K}(\mathbb{T})\oplus K$ is a Type II invariant subspace for $B_{\sigma,e^{i\theta}}^{K}$, and so will be
$H^{2}_{\{0\}\oplus G}(\mathbb{T})\oplus(\{0\}\oplus G)=\left(H^{2}_{\{0\}\oplus G}(\mathbb{T})\oplus\{0\}\right)\oplus\left(\{0\}\oplus(\{0\}\oplus G)\right)$
for $B_{\sigma^{\prime},e^{i\theta^{\prime}}}^{K}$. However, in view of (6.2), condition (2) of Theorem 3.1 guarantees the existence of unitary operators $U^{\prime}_{E}$ and $U^{\prime}_{E_{1}}$ such that
$U^{\prime}_{E}(E\oplus\{0\})=\{0\}\oplus G$
and
$U^{\prime}_{E_{1}}(\{0\}\oplus(E_{1}\oplus\{0\}))=\{0\}\oplus(\{0\}\oplus G).$
Clearly, these two unitary operators $U^{\prime}_{E}$ and $U^{\prime}_{E_{1}}$ induce two other unitary operators $U_{E}$ and $U_{E_{1}}$ such that
$U_{E}(E)=G\text{ and }U_{E_{1}}(E_{1})=G.$
Hence, $U_{E_{1}}^{*}U_{E}$ defines a unitary operator between $E$ and $E_{1}$, leading to the conclusion:
$\dim E=\dim E_{1},$
which is contrary to our assumption. Therefore, the operator $B_{\sigma,e^{i\theta}}^{E}\big{|}_{H^{2}_{E}(\mathbb{T})\oplus E_{1}}$ is never unitarily equivalent to any other Brownian shift.
In the following example, we show that the restriction of a Brownian shift to an invariant subspace may result in operator that is unitarily equivalent to another Brownian shift.
Example 6.2.
Consider the Brownian shift $B_{\sigma,1}^{\mathbb{C}^{2}}$ on $H^{2}_{\mathbb{C}^{2}}(\mathbb{T})\oplus\mathbb{C}^{2}$. We use the row-vector notation $\begin{bmatrix}f_{1}&f_{2}\end{bmatrix}$, where $f_{1},f_{2}\in H^{2}(\mathbb{T})$, to denote an element of $H^{2}_{\mathbb{C}^{2}}(\mathbb{T})$. Note that, in particular, either $f_{1}$ or $f_{2}$ may be a constant in $\mathbb{C}$. However, for $\begin{bmatrix}0&0\end{bmatrix}$, we will simply write 0 when there is no chance of confusion. Now, we begin by considering the one-dimensional subspace of $\mathbb{C}^{2}$ as
$E_{1}=\langle[1,1]\rangle.$
Also, consider the inner function $\Phi\in H^{\infty}_{\mathcal{B}(E_{1},\mathbb{C}^{2})}(\mathbb{T})$, defined by $\Phi(z)=\Phi_{1}(z)\big{|}_{E_{1}}$, where
$\Phi_{1}(z)=\begin{bmatrix}z&0\\ 0&1\end{bmatrix}$.
Note that $\Phi_{1}\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{2})}(\mathbb{T})$ is an inner function. Since we can write $\begin{bmatrix}\sigma&0\end{bmatrix}$ as
$\begin{bmatrix}\sigma&0\end{bmatrix}=\frac{\Phi\begin{bmatrix}\sigma&\sigma\end{bmatrix}-\sigma\begin{bmatrix}1&1\end{bmatrix}}{z-1}\in H^{2}_{\mathbb{C}^{2}}(\mathbb{T}),$
by recalling the definition of $\mathcal{G}_{\Phi}$ from (1.1), it follows that
$\mathbb{g}=\begin{bmatrix}\begin{bmatrix}\sigma&0\end{bmatrix}\\ \\ \begin{bmatrix}1&1\end{bmatrix}\end{bmatrix}$
is a member of $\mathcal{G}_{\Phi}$. Therefore, Theorem 2.5 ensures that
$\mathcal{M}=(\Phi H^{2}_{E_{1}}(\mathbb{T})\oplus\{0\})\oplus\langle\mathbb{g}\rangle,$
is a Type II invariant subspace of $B_{\sigma,1}^{\mathbb{C}^{2}}$. At this point, we turn our attention to the Brownian shift
$B_{\sigma^{\prime},1}^{\mathbb{C}\oplus\{0\}}\in\mathcal{B}\left(H^{2}_{\mathbb{C}\oplus\{0\}}(\mathbb{T})\oplus\left(\mathbb{C}\oplus\{0\}\right)\right),$
for some $\sigma^{\prime}>0$. We claim that
$B_{\sigma,1}^{\mathbb{C}^{2}}\big{|}_{\mathcal{M}}\cong B_{\sigma^{\prime},1}^{\mathbb{C}\oplus\{0\}}=B_{\sigma^{\prime},1}^{\mathbb{C}^{2}}\big{|}_{H^{2}_{\mathbb{C}\oplus\{0\}}(\mathbb{T})\oplus\left(\mathbb{C}\oplus\{0\}\right)},$ (6.3)
where
$\sigma^{\prime}=\frac{\sigma\sqrt{2}}{\sqrt{2+\sigma^{2}}}>0.$
It is evident that $H^{2}_{\mathbb{C}\oplus\{0\}}(\mathbb{T})\oplus\left(\mathbb{C}\oplus\{0\}\right)$ is a Type II invariant subspace of $B_{\sigma^{\prime},1}^{\mathbb{C}^{2}}\in\mathcal{B}(H^{2}_{\mathbb{C}^{2}}(\mathbb{T})\oplus\mathbb{C}^{2})$, if we recognize that
$H^{2}_{\mathbb{C}\oplus\{0\}}(\mathbb{T})\oplus\left(\mathbb{C}\oplus\{0\}\right)=\left(H^{2}_{\left\langle\begin{bmatrix}1&0\end{bmatrix}\right\rangle}(\mathbb{T})\oplus\{0\}\right)\oplus\mathbb{C}\begin{bmatrix}\begin{bmatrix}0&0\end{bmatrix}\\ \\ \begin{bmatrix}1&0\end{bmatrix}\end{bmatrix}$,
and that
$\begin{bmatrix}0&0\end{bmatrix}=\frac{\begin{bmatrix}\sigma^{\prime}&0\end{bmatrix}-\sigma^{\prime}\begin{bmatrix}1&0\end{bmatrix}}{z-1}.$
With all the above information, we are now able to define a unitary $U_{\mathbb{g}}$ by
$U_{\mathbb{g}}(\mathbb{g})=\sqrt{2+\sigma^{2}}\begin{bmatrix}\begin{bmatrix}0&0\end{bmatrix}\\ \\ \begin{bmatrix}1&0\end{bmatrix}\end{bmatrix}$,
and another unitary $U_{E_{1}}$ by
$U_{E_{1}}\left(\begin{bmatrix}\sigma&\sigma\end{bmatrix}\right)=\sqrt{2+\sigma^{2}}\begin{bmatrix}\sigma^{\prime}&0\end{bmatrix}$.
The condition (2) of Theorem 3.1 now guarantees that (6.3) is true.
In light of the preceding two examples, it is now clear that Brownian shifts have distinct characteristics that are not commonly observed in shifts on Hilbert function spaces.